# Semi-supervised Anomaly Detection via Adaptive Reinforcement Learning-Enabled Method with Causal Inference

### v1

@@ 4.3 Causal SAC
To find the essential causal feature using CI and raw data, we need to model the relationship between them. We divide this modeling process into two categories: intra-point relationship modeling and inter-point relationship modeling. Intra-point relationship modeling refers to the use of the graph attention mechanism to embed the dimensions of data points into the attention graph, to discover the relationship between inner dimensions. Inter-point relationship modeling refers to the use of the use of CI to remove the confounding factors in the original data to extract the intrinsic causal feature. As shown in the causal graph (Figure 2), $N$ is the confounding factor that needs to be removed. First, we use the backdoor formula as shown in Equation (2) for this causal inference process. Because the counterfactual data $c$ required for CI is unknown, to utilize the counterfactual data for intervention ($\text{do}(X=c)$), we replace each feature dimension in the original data with the mean of all dimensions as the counterfactual data. As described in Section 3.3, the counterfactual data intervene in the causal graph through the backdoor formula, which can make the path $N\rightarrow F$ be blocked, avoiding the unfavorable influence of $N$ on $L$. Then, to obtain the intrinsic causal feature $F$, we construct the Causal Feature Extractor (CFE) module, which models the feature information of the original data with that of the counterfactual intervention and explores the causal relationship between the two features, to infer the intrinsic causal feature $F$ associated with each data point. The process is shown in the following equation:
@@ 5 Experiments
This section will present the dataset used in this work, the evaluation metrics, and the results of Tri-CRLAD compared to 9 baseline methods.
@@ 5.1 Datasets
As shown in Table 1, to comprehensively assess the generalizability of the Tri-CRLAD, four single-anomaly datasets and three multiple-anomaly datasets were selected for the study. These datasets cover different application domains ranging from medical diagnosis to satellite monitoring, ensuring the diversity and breadth of the study.
The single anomaly datasets include annthyroid, cardio, satellite, and satimage. The percentage of anomaly data in these datasets ranges from 1.2% to 31.6%. These datasets were chosen to demonstrate the model’s effectiveness in dealing with single-anomaly scenarios with varying proportions of anomalies. On the other hand, Multi-anomaly datasets include Multi-annthyroid for disease diagnosis, cardio for fetal cardiotocogram diagnosis, and har for human activity recognition. The proportion of single-category anomalies in these datasets ranges from 2.5% to 15%. These diverse datasets with multiple types of abnormalities help evaluate the model’s performance in dealing with complex scenarios.
@@ 5.4 Baselines and Comparisons
We have chosen a few methods, including supervised, unsupervised, and semi-supervised learning, as our baseline methods for this experiment. These methods include the popular XGBoost [37] and Isolation Forest [38], as well as two methods in the same field as our work - DADS [7] and DPLAN [8]. Additionally, we have selected several other SOTA methods for semi-supervised learning, such as DeepSAD [39], DevNet [21], SSAD [40] and STOC [18] method. Notably, we cite the experimental results in DADS and the STOC+GOAD [41] set up in that paper with the results of STOC+IForest as the performance of STOC in this experiment. The experimental results for DeepSAD correspond to the performance of an upgraded version of this model (Hypersphere Classification (HSC) [42]).
In this experimental section, we will show the robustness of the proposed model step by step. First, we tested the model’s ability to perform in a single anomaly dataset under an anomalies ratio of 0.1 and a contamination ratio of 0.1. This part of the experiment aims to test the fundamental ability that should be possessed as an anomaly detection model, i.e., the ability to detect anomaly data points accurately. As demonstrated in Table LABEL:tab2, Tri-CRLAD excels at utilizing a minimal number of labeled anomalies to effectively identify unlabeled anomaly data points within the $D_{u}$.
@@
Next, we further tested Tri-CRLAD’s ability to perform in a multi-anomaly scenario. This scenario implies that an agent can detect anomalies containing a known class and other classes from unlabeled data when only a very small number of labeled anomalies within a class of anomalies are known, which is undoubtedly challenging. This part of the experiment can be divided into two parts, one in which the anomalies ratio is fixed and the contamination ratio is varied (Scenario 1), and the other in which the contamination ratio is fixed and the anomalies ratio is varied (Scenario 2). In Scenario 1, we fixed the anomaly ratio at 0.1 and adjusted the contamination ratio in increments of 0.02 within a range of [0, 0.1]. Figure 4 shows the model’s performance with different datasets, and Table LABEL:tab3 exhibits the results when the anomalies ratio=0.1 and the contamination ratio=0.1. The outcomes of these experiments reveal that the model demonstrates considerable stability when facing gradual changes in contamination levels. This finding affirms the model’s robustness and its capability to adapt effectively in environments characterized by multiple and challenging anomaly detection tasks within data.
@@
In Scenario 2, this part of the experiment fixes the contamination ratio at 0.04, and the anomalies ratio is varied in the range of 0.01, 0.05, 0.1, 0.15, and 0.5. The results presented in Table 4 show the performance of the model for anomalies ratio=0.01, and Figure 5 shows the performance of each model corresponding to each anomalies ratio. In the condition of anomalies ratio=0.01, the available prior knowledge is greatly reduced, and the capabilities of the various models show a significant decline. However, due to the effective collaboration between the CFE module and the triple decision support mechanism within ADIE, Tri-CRLAD exhibits resilience in this challenging scenario. Its maximum recession rate is limited to 15% compared to the peak performance observed in this experimental segment. In contrast, other methods experienced a much steeper decline, with maximum recession rates reaching as high as 38%. Consequently, considering the overall results, Tri-CRLAD maintains remarkable stability without significant fluctuations, further underscoring the model’s exceptional robustness.
@@
Compared to other semi-supervised or unsupervised algorithms for non-reinforcement learning, Tri-CRLAD demonstrates a clear lead by its superior ability to utilize prior knowledge. Overall, Tri-CRLAD’s success is attributed to its efficient sampling strategy, which motivates the agent to explore more extensively. Next, the intrinsic causal feature extracted through causal inference, combined with the adaptive decision reward mechanism, provides comprehensive learning information for the agent. Finally, with the help of the adaptive smoothing threshold adjustment strategy, the model can comprehensively perceive the learning situation to determine the abnormal data accurately.
@@ 5.5 Ablation study
First, we analyze the role of the adaptive threshold smoothing strategy. The adaptive threshold smoothing strategy is designed to enable the agent to determine the abnormal data more accurately. As can be seen from the results in Table 5, compared with the fixed thresholds, the adaptive changing thresholds enable the agent to detect anomalies more accurately. Meanwhile, to further demonstrate the superiority of the structure described in the method section, Figure 6 demonstrates the difference in the changing trend under the non-adaptive adjustment strategy and the adaptive adjustment strategy constructed in this paper, and Table 5 also contains the impact of the two different results on the performance of Tri-CRLAD. This series of sufficient experiments shows that the adaptive threshold smoothing strategy constructed in this paper makes the adjustment of the threshold smoother on the one hand, making the model’s decision more stable. On the other hand, the strategy improves the detection accuracy of the model, which can be better suited for various tasks.
@@ 6 Conclusion
In this study, we propose the Tri-CRLAD model, a new approach to address the key challenges facing semi-supervised anomaly detection in the field of reinforcement learning. The model introduces causal reinforcement learning into semi-supervised anomaly detection and solves the problem of underutilizing prior knowledge by mining the intrinsic causal properties of data through counterfactual reasoning, which improves the ability of the agent to utilize prior knowledge. In addition, the model proposes a triple decision support mechanism based on historical information, which enhances the model’s flexibility and improves the problem of insufficient reward feedback when interacting with the environment. Through these innovations, Tri-CRLAD can effectively respond to multiple complex and changing environments, demonstrating a wide range of potential applications. Ultimately, Tri-CRLAD demonstrates superior performance on anomaly detection datasets across multiple domains, meeting or exceeding existing state-of-the-art semi-supervised anomaly detection methods.

### v2

@@ 4.3 Causal SAC
To find the essential causal feature using CI and raw sensor signals, we need to model the relationship between them. We divide this modeling process into two categories: intra-point relationship modeling and inter-point relationship modeling. Intra-point relationship modeling refers to the use of the graph attention mechanism to embed the dimensions of data points into the attention graph, to discover the relationship between inner dimensions. Inter-point relationship modeling refers to the use of the use of CI to remove the confounding factors in the original data to extract the intrinsic causal feature. As shown in the causal graph (Figure 2), $N$ is the confounding factor that needs to be removed. First, we use the backdoor formula as shown in Equation (2) for this causal inference process. Because the counterfactual data $c$ required for CI is unknown, to utilize the counterfactual data for intervention ($\text{do}(X=c)$), we replace each feature dimension in the original data with the mean of all dimensions as the counterfactual data. As described in Section 3.3, the counterfactual data intervene in the causal graph through the backdoor formula, which can make the path $N\rightarrow F$ be blocked, avoiding the unfavorable influence of $N$ on $L$. Then, to obtain the intrinsic causal feature $F$, we construct the Causal Feature Extractor (CFE) module, which models the feature information of the original data with that of the counterfactual intervention and explores the causal relationship between the two features, to infer the intrinsic causal feature $F$ associated with each data point. The process is shown in the following equation:
@@ 5 Experiments
This section will present the sensor signals dataset used in this work, the evaluation metrics, and the results of Tri-CRLAD compared to 9 baseline methods.
@@ 5.1 Datasets
To comprehensively assess the generalizability of the Tri-CRLAD, we selected a diverse set of single-anomaly and multiple-anomaly datasets in the field of sensor signals, as shown in Table 1. These datasets cover various application areas, including medical monitoring, health diagnostics, behavior monitoring, and satellite monitoring, ensuring a broad and diverse scope for our research.
The single-anomaly datasets—annthyroid, cardio, satellite, and satimage—feature anomaly data percentages ranging from 1.2% to 31.6%. These datasets were specifically chosen to evaluate the model’s effectiveness in handling single-anomaly scenarios with varying proportions of anomalies. On the other hand, the multiple-anomaly datasets include Multi-annthyroid for disease diagnosis, cardio for fetal cardiotocogram diagnosis, and har for human activity recognition. The proportion of single-category anomalies in these datasets ranges from 2.5% to 15%. These diverse datasets, encompassing multiple types of abnormalities, provide a robust evaluation of the model’s performance in complex scenarios involving sensor signals. Utilizing these sensor data enables the diagnosis of thyroid disease and the monitoring of cardiovascular health.
@@ 5.4 Baselines and Comparisons
We have chosen various methods, including supervised, unsupervised, and semi-supervised learning, as our baseline methods for this experiment. These methods include the popular XGBoost [38] and Isolation Forest [39], as well as two methods in the same field as our work - DADS [8] and DPLAN [9]. Additionally, we have selected several other SOTA methods for semi-supervised learning, such as DeepSAD [40], DevNet [23], SSAD [41] and STOC [20] method. Notably, we cite the experimental results in DADS and the STOC+GOAD [42] set up in that paper with the results of STOC+IForest as the performance of STOC in this experiment. The experimental results for DeepSAD correspond to the performance of an upgraded version of this model (Hypersphere Classification (HSC) [43]).
In this experimental section, we will show the robustness of the proposed model step by step. First, we tested the model’s ability to perform in a single anomaly dataset under an anomalies ratio of 0.1 and a contamination ratio of 0.1. This part of the experiment aims to test the fundamental ability that should be possessed as an anomaly detection model, i.e., the ability to detect anomaly sensor signal data points accurately. As demonstrated in Table LABEL:tab2, Tri-CRLAD excels at utilizing a minimal number of labeled anomalies to effectively identify unlabeled anomaly data points within the $D_{u}$.
@@
Next, we further tested Tri-CRLAD’s ability to perform in a multi-anomaly scenario. This scenario implies that an agent can detect anomalies containing a known class and other classes from unlabeled data when only a very small number of labeled anomalies within a class of anomalies are known, which is undoubtedly challenging. This part of the experiment can be divided into two parts, one in which the anomalies ratio is fixed and the contamination ratio is varied (Scenario 1), and the other in which the contamination ratio is fixed and the anomalies ratio is varied (Scenario 2). In Scenario 1, we fixed the anomaly ratio at 0.1 and adjusted the contamination ratio in increments of 0.02 within a range of [0, 0.1]. Figure 4 shows the model’s performance with different datasets, and Table LABEL:tab3 exhibits the results when the anomalies ratio=0.1 and the contamination ratio=0.1. The outcomes of these experiments reveal that the model demonstrates considerable stability when facing gradual changes in contamination levels. This finding affirms the model’s robustness and its capability to adapt effectively in environments characterized by multiple and challenging anomaly detection tasks within sensor signal data.
@@
In Scenario 2, this part of the experiment fixes the contamination ratio at 0.04, and the anomalies ratio is varied in the range of 0.01, 0.05, 0.1, 0.15, and 0.5. The results presented in Table 4 show the performance of the model for anomalies ratio=0.01, and Figure 5 shows the performance of each model corresponding to each anomalies ratio. In the condition of anomalies ratio=0.01, the available prior knowledge is greatly reduced, and the capabilities of the various models show a significant decline. However, due to the effective collaboration between the CFE module and the triple decision support mechanism within ADIE, Tri-CRLAD exhibits resilience in this challenging scenario. Its maximum recession rate is limited to 15% compared to the peak performance observed in this experimental segment. In contrast, other methods experienced a much steeper decline, with maximum recession rates reaching as high as 38%. Consequently, considering the overall results, Tri-CRLAD maintains remarkable stability without significant fluctuations, further underscoring the model’s exceptional robustness for dealing with sensor signals.
@@
Compared to other semi-supervised or unsupervised algorithms for non-reinforcement learning, Tri-CRLAD demonstrates a clear lead by its superior ability to utilize prior knowledge. Overall, Tri-CRLAD’s success is attributed to its efficient sampling strategy, which motivates the agent to explore more extensively. Next, the intrinsic causal feature extracted through causal inference, combined with the adaptive decision reward mechanism, provides comprehensive learning information for the agent. Finally, with the help of the adaptive smoothing threshold adjustment strategy, the model can comprehensively perceive the learning situation to determine the abnormal sensor signal data accurately.
@@ 5.5 Ablation Study
First, we analyze the role of the adaptive threshold smoothing strategy. The adaptive threshold smoothing strategy is designed to enable the agent to determine the abnormal data more accurately. As can be seen from the results in Table 5, compared with the fixed thresholds, the adaptive changing thresholds enable the agent to detect anomalies more accurately. Meanwhile, to further demonstrate the superiority of the structure described in the method section, Figure 6 demonstrates the difference in the changing trend under the non-adaptive adjustment strategy and the adaptive adjustment strategy constructed in this paper, and Table 5 also contains the impact of the two different results on the performance of Tri-CRLAD. This series of sufficient experiments shows that the adaptive threshold smoothing strategy constructed in this paper makes the adjustment of the threshold smoother on the one hand, making the model’s decision more stable. On the other hand, the strategy improves the detection accuracy of the model, which can be better suited for various sensor signal tasks.
@@ 6 Conclusion
In the era of smart manufacturing, enhanced sensing and computing capabilities have made anomaly detection essential in various sensor-based applications. In this study, we propose the Tri-CRLAD model, a new approach to address the key challenges facing semi-supervised anomaly detection in the field of reinforcement learning. The model introduces causal reinforcement learning into semi-supervised anomaly detection and solves the problem of underutilizing prior knowledge by mining the intrinsic causal properties of data through counterfactual reasoning, which improves the ability of the agent to utilize prior knowledge. In addition, the model proposes a triple decision support mechanism based on historical information, which enhances the model’s flexibility and improves the problem of insufficient reward feedback when interacting with the environment. Through these innovations, Tri-CRLAD can effectively respond to complex and changing sensor signal environments, demonstrating a wide range of potential applications. Ultimately, Tri-CRLAD demonstrates superior performance on multiple anomaly detection datasets in the sensor signal domain, meeting or exceeding existing state-of-the-art semi-supervised anomaly detection methods.